# Tensor Networks for Explainable Machine Learning in Cybersecurity

### v3

@@ II Methodology
This section delves into the adequacy of MPS within the scope of unsupervised generative modeling for anomaly detection. We elucidate the mechanisms for extracting valuable insights from MPS, particularly through the concept of reduced density matrices, proving the versatility of MPS in modeling complex data distributions and distilling meaningful information from its structure.
II.1 Unsupervised generative modeling with MPS
@@
II.2 Extracting information from MPS
@@ III Adversary-Generated threat intelligence
For this investigation, we delve into a dataset of cyber-attack vectors obtained from real-life scenarios, and provided to us by the cybersecurity company CounterCraft SL. Encompassing a variety of configurations and system states, this dataset mirrors the intricate nuances and complexities that characterize true cyber-attacks. It is against this collection of diverse cyber threats that we will later unpack the interpretability brought up by MPS.
@@
More specifically, the dataset deployed in this study is a collection of five distinct subsets, each corresponding to the totality of events recorded across all systems on a specific day. This compilation method offers a longitudinal view of cyber activities, gathering both benign and malicious events. As evidenced by the data, there is a pronounced imbalance between the overall number of events and those identified as part of cyber-attack incidents, which is reflective of the asymmetry typically encountered in real-world cybersecurity datasets. The subsets, named here ‘CC1’ through ‘CC5’, present a diverse landscape of interactions, with ’CC1’ containing $140,202$ events of which 34 are incident-related, scaling up to ‘CC5’ with $194,766$ events including 822 tied to reported incidents. In total, we have $674,704$ events of which $1,007$ are incident-related. This realistic imbalance poses a unique challenge, and underscores the need for a robust anomaly detection system capable of discerning subtle patterns of cyber-threats within predominantly benign data traffic.
@@
Our attack-detection system will identify attacks as anomalies. Notably, not all detected anomalies will be attacks. Indeed, there may be low-probability events that are perfectly fine (say, energy outages, reboots of the computer systems…). Because of this, it is important to play with the different parameters available in the model to minimize the detection of false positives, i.e., anomalies that are not attacks after all. With this in mind, central to the model’s success is the implementation of an acceptance threshold for the Negative Log Likelihood (NLL) of each event being generated by the model, so that events that exceed the threshold are classified as anomalous and may be considered by a human analyst to check if they correspond to a real attack. Establishing this threshold is challenging and varies in practice; some analysts may use statistical metrics like the mean and standard deviation to discern attacks within anomalies, while others may opt for the median. Through rigorous training, which leverages 70% of the data and respects the chronological order of events, our model effectively captures attack patterns. This temporal fidelity is crucial as shuffling the order of events has been shown to influence model training negatively. Notably, the average NLL of daily benign events is consistently lower than that of attacks, confirming the model’s adeptness at flagging attacks as anomalies.
@@
In Fig.3 we show one way to determine the threshold of NLL in our model, by analyzing the number of anomalies detected and the number of attack’s events present in those detected anomalies for multiple threshold values (notice that, in the figure, the scales for both quantities are different). Depending on the specific goal, one can choose a low threshold value and maximise the number of attack’s events at the expense of having a higher false positive rate, or choose a high value for the threshold and minimise the false positive rate at the expense of not detecting all the attack’s events inside the detected anomalies. In any case, in the figure we can see that the number of detected anomalies decreases very quickly with the threshold, whereas this is not the case for the number of detected attacks, meaning that our MPS model is remarkably good at pinpointing anomalies that correspond to true attacks.
In practice, our MPS model achieves remarkable results: we see that by fine-tuning the thershold, an average of 83.5% of incident-report events are correctly identified as anomalies, and all attack types are detected (100% attack detection rate). Moreover, the average false positive rate (FPR) stands at a manageable 1.39%, substantially narrowing the search space for analysts by 98.61%. As shown in Fig. 4, the FPR decreases when increasing the virtual bond dimension of the TN, outperforming in this metric many traditional machine learning methods and comparing to that of autoencoders. These FPR values are achieved while maintaining a high detection precision of 92.3%. Notably, the results could be improved even further by increasing the MPS bond dimension, and/or by using more complex TN structures, able to capture more subtle correlations in the model. Last but not least, our MPS model can be scaled easily to datasets with a huge number of features, and is remarkably memory-efficient, on top of being accurate (as shown) and explainable. This last property is what we discuss in detail in the next section.
@@ IV.1 Direct probability extraction
As presented in subsection II.2, the RDM’s diagonal entries provide a probabilistic view of specific subsystem’s state, which is directly connected to the dataset’s frequency distribution. The MPS captures and adapts to the correlations in the data during its training phase, and the RDM can reflect then the empirical frequency distribution that takes into account the correlations between the different features. While a simple frequency distribution is akin to a static snapshot, capturing conditional probabilities at a fixed time step, MPS allows for an efficient dynamical representation of other valuable insights, turning exhaustive analysis such as the case of all-to-all feature correlations (scaling exponentially in computational complexity) into a polynomial complexity scaling with memory requirement scaling as $O(D^{3})$, with $D$ the bond dimension.
@@ IV.2 Von Neumann entropy
The concept of Von Neumann Entropy is analogous to the Shannon entropy in classical information theory [16] but adapted for quantum systems and reduced density matrices. In the end, the Von Neuamn entropy of a RDM is nothing but the Shannon entropy of its eigenvalues. It quantifies the correlation between a subsystem with the rest of the system, thus providing insight into how independent a feature is within the broader data context. This metric is pivotal in our analysis of anomaly detection, and provides insight into which features are independent and which are context-dependent. This allows to differentiate between what might be deemed normal and anomalous within a dataset, effectively utilizing “context” to pinpoint deviations.
@@ IV.3 Feature importance
  Feature Index   Benign   Attack
@@
Total: 2.14e-10 6.06e-18
Table 1: Mean probabilities of feature values for benign and attack instances. The total is the product of mean probabilities across all features, reflecting the overall likelihood of benign versus attack profiles.
Table 1 presents the computed mean probabilities for features in benign and attack contexts, which can be instrumental in feature selection. Features with high probabilities in the benign category and low probabilities in the attack category, such as features 2, 3, and 4, are indicative of normal operational patterns. Their high likelihood under benign conditions and rarity during attacks signify their importance in identifying abnormal behavior. Conversely, features with a higher mean in the attack category than in the benign may be less informative for detecting anomalies, as their higher occurrence in attack scenarios does not contribute to distinguishing attacks from benign behavior. In addition, the total probability values, obtained by the product of probabilities across all features, reveal a striking disparity: the probability associated with attack features is several orders of magnitude lower than that of benign features. This stark contrast explains why the MPS model effectively discriminates between benign operations and cyberattacks: the improbability of the conjunction of attack features underscores their anomaly.
@@ IV.6 Mutual Information
The incorporation of mutual information into our MPS framework allows to understand the intricate correlations between features within the system. The mutual information heatmap depicted in Fig.9 provides a vivid portrayal of feature interdependencies, offering a window into the dataset’s complex interrelations. This analytical tool is invaluable for pinpointing pairs of features with significant shared information, thereby enhancing our grasp of the system’s underlying architecture. Beyond interpretability, this metric holds practical utility in optimizing the MPS model. By rearranging columns to position highly correlated features adjacent to each other, we can refine the efficiency of the virtual bond dimension, ensuring a more resource-effective and potent representation of the data’s entangled nature. In addition, those pairs of features with large mutual information could even be clustered together, therefore reducing the complexity of the problem.
@@ V Conclusions and Next Steps
In this paper we have discussed in detail the potential of TN methods to develop explainable and interpretable machine learning algorithms. We have shown the validity of the approach for an MPS algorithm designed for clustering, with the application to a real-life use case in cybersecurity. Our research has unveiled the potential of MPS as a powerful tool for unsupervised anomaly detection, particularly within the domain of adversary-generated threat intelligence. The key findings demonstrate that MPS not only rivals but in certain respects surpasses the performance of traditional machine learning models such as autoencoders and GANs, especially in terms of explainability and interpretability. The application of MPS to a cybersecurity dataset revealed its robustness in identifying cyberthreats with a low false positive rate, a critical parameter in cybersecurity operations. The ability of MPS to provide direct probability extraction, assess feature importance, and elucidate on the reasons behind the classification of data points as anomalies, underscores its utility in operational environments where understanding the ‘why’ behind a model’s decision is as vital as the decision itself. In addition, the research has highlighted the interpretability capabilities of MPS, offering insights into Von Neumann entropy, feature importance, conditional probabilities, and mutual information directly from the model’s parameters without the need for additional tools or techniques. These capabilities have profound implications for explainable AI, setting a new benchmark for transparency in machine learning applications. In addition, the methods presented in this paper allow in the capability to generate synthetic data in order to improve training of complex models (due to lack of data availability) and activity generation for deception systems. This would help also in the development of behavior-based models, which are better against unknown attacks than rule-based models.
The methods developed in this work can be improved in many ways. For instace, we could consider more advanced TN structures beyond MPS, with more intrincate patterns of correlations. We leave this for future studies.
@@
Acknowldgements.- We acknowledge Donostia International Physics Center (DIPC), Ikerbasque, Basque Government, Diputación de Gipuzkoa, European Innovation Council (EIC) and Tecnun for constant support, as well as insightful discussions with the teams from Multiverse Computing, DIPC and Tecnun on the algorithms and technical implementations. Special thanks to the team at CounterCraft SL, for providing the motivation for this work, the dataset and numerous technical discussions.

### v4

@@ II Methodology
This section explores the role of TNs in explainable machine learning, with a focus on MPS as a powerful framework for anomaly detection. We first introduce the general concept of TNs and their relevance to high-dimensional data representations. Then, we delve into the adequacy of MPS within the scope of unsupervised generative modeling for anomaly detection, illustrating its ability to efficiently capture feature correlations. We also explain the mechanisms for extracting valuable insights from MPS, particularly through the use of reduced density matrices, which provide deeper interpretability in machine learning applications.
II.1 Introduction to Tensor Networks
Tensor Networks (TNs) are a mathematical framework designed to efficiently represent and manipulate high-dimensional data. They originate from quantum many-body physics and have found applications in machine learning, classical statistical mechanics, quantum chemistry, and even finance [13, 12, 14].
The central idea of TNs is to decompose high-dimensional tensors into a network of lower-dimensional tensors connected by shared indices. This structured representation helps in efficiently capturing correlations in the data, reducing computational complexity, and maintaining interpretability.
Several types of TN architectures exist, each suited for different types of problems [6, 15]:
  • Matrix Product States (MPS): The simplest and most commonly used TN, ideal for one-dimensional sequences, as in anomaly detection.
  • Projected Entangled Pair States (PEPS): A two-dimensional generalization of MPS, used for more complex correlations but computationally demanding.
  • Multiscale Entanglement Renormalization Ansatz (MERA): Suitable for hierarchical data and quantum simulations, leveraging scale invariance.
In the context of machine learning, TNs enable structured compression of data while preserving essential relationships between features. This ability makes them particularly attractive for interpretable AI applications. Among these TN architectures, MPS stands out as a powerful tool for anomaly detection, efficiently capturing correlations between sequential features while remaining computationally tractable. In the next subsection, we detail the use of MPS in our approach.
II.2 Unsupervised generative modeling with MPS
$P(v)=\left|\Psi\left(v\right)\right|^{2}$ (1)
where $\Psi(v)$ is the coefficient of the normalized quantum state
$|\Psi\rangle=\sum_{v\in V}\Psi\left(v\right)|v\rangle,$ (2)
where each $A^{\left(k\right)v_{k}}\,$ is a $D_{k-1}\times D_{k}\,$ matrix.
II.3 Unsupervised generative modeling with MPS
@@
The choice of bond dimension D directly affects the computational complexity of the MPS model, scales as $O(NpD^{3})$ with $N$ number of features and $p$ physical dimension. In our implementation, we find that increasing D beyond a certain threshold ($\sim$ 150) leads to diminishing returns in anomaly detection performance while significantly increasing training time.
@@
II.4 Extracting information from MPS
@@ III Adversary-Generated Threat Intelligence
For this investigation, we delve into a dataset of cyber-attack vectors obtained from real-life scenarios, provided to us by the cybersecurity company CounterCraft SL. Encompassing a variety of configurations and system states, this dataset mirrors the intricate nuances and complexities that characterize true cyber-attacks. It is against this collection of diverse cyber threats that we later unpack the interpretability brought up by MPS.
@@
More specifically, the dataset deployed in this study is a collection of five distinct subsets, each corresponding to the totality of events recorded across all systems on a specific day. This compilation method offers a longitudinal view of cyber activities, gathering both benign and malicious events. As evidenced by the data, there is a pronounced imbalance between the overall number of events and those identified as part of cyber-attack incidents, which is reflective of the asymmetry typically encountered in real-world cybersecurity datasets. The subsets, named here ‘CC1’ through ‘CC5’, present a diverse landscape of interactions, with ‘CC1’ containing $140,202$ events of which 34 are incident-related, scaling up to ‘CC5’ with $194,766$ events including 822 tied to reported incidents. In total, we have $674,704$ events of which $1,007$ are incident-related. This realistic imbalance poses a unique challenge and underscores the need for a robust anomaly detection system capable of discerning subtle patterns of cyber-threats within predominantly benign data traffic.
To further elucidate the structure and complexity of the dataset, we provide a snapshot of sample entries in Table 1. Each entry represents a system event characterized by multiple features, including process paths, usernames, event types, and outcomes:
Table 1: Sample entries from the dataset highlighting key features.
Parent Process Path Process User Parent Process Process Name Real User Process Path Type Code Category Outcome Directory Host IP /usr/sbin /nginx www-data dash hostname root /usr/bin /sed Connection Established process UNKNOWN /usr/bin 10.0.0.4 /snap/amazon-ssm-agent/6312/ssm-agent-worker root ssm-agent-worker sed www-data /usr/sbin /nginx Create Process network SUCCESS /usr/sbin 172.31.3.38 /usr/bin /dash root nginx nginx root /usr/bin /hostname DNSQuery network SUCCESS /usr/bin 172.31.14.43
The table highlights critical features such as:
  • Process Information: Details about the parent and child processes (‘Parent Process Path‘, ‘Parent Process‘, ‘Process Path‘, ‘Process Name‘).
  • User Details: The user responsible for the processes (‘Process User‘, ‘Real User‘).
  • Event Metadata: Event type (‘Type Code‘), associated tags (‘Tags‘), category (‘Category‘), and outcome (‘Outcome‘).
  • Network Data: Host IP address and directories involved.
These features encapsulate the breadth of system activities, enabling the detection of cyber-attacks as anomalies. In addition to the core features, an extra Tag feature is derived to classify events based on the MITRE ATT&CK framework and proprietary classifications such as CounterCraft’s unique TTPs (e.g., T9xxx series). Events tagged with specific threat indicators signal potential attacks, such as attack::T9007.002, providing critical insights for identifying malicious activities. The diversity and granularity of these features are essential for the MPS model to effectively discern malicious activities while maintaining interpretability. In the following figures, a total of 23 features are presented. These features were derived by applying numerical encoding and distribution techniques to the original feature set described above, allowing for enhanced representation and facilitating more effective MPS training.
Our methodology for applying the MPS model to anomaly detection in adversary-generated threat intelligence consists of the following steps:
  1. 1.
    Data Preprocessing:
    • Extract features from logs in various formats (JSON, CSV, PCAP).
    • Filter out numeric features with zero variance (constant values) and remove duplicated features.
    • Perform correlation analysis (Spearman and Cramér’s V) to remove highly correlated numeric features.
    • If labels are available, analyze feature correlation with labels and select the most important features.
    • Apply clustering techniques and visualization for exploratory analysis.
    • Encode categorical features using one-hot encoding or numerical encoding.
    • Perform time dependency analysis to check for data leakage.
    • Normalize the dataset and reduce the number of features using Principal Component Analysis (PCA).
  2. 2.
    Feature Encoding:
    • The selected features are numerically encoded to align with the MPS model’s input format. This ensures compatibility with the tensor network structure and optimizes the representation of both categorical and continuous data.
  3. 3.
    Training the MPS Model:
    • The MPS model is trained on the first 70% of benign events, sorted chronologically to ensure a realistic baseline.
    • The MPS model learns the statistical distribution of normal behavior, capturing feature correlations.
    • Hyperparameter tuning is performed to optimize performance, balancing computational efficiency and accuracy. The key hyperparameters include:
      • MaxBondDim: Maximum bond dimension of the MPS, controlling the model’s expressiveness.
      • lr_shrink: Learning rate reduction factor if the loss stops improving.
      • loopmax: Maximum number of training loops.
      • safe_thres: Threshold for detecting unstable learning rates.
      • lr_inf: Minimum learning rate before training stops.
      • lr_var: Initial learning rate.
      • dcs: Number of descent steps within each optimization loop.
      • cut: Threshold for truncating small singular values in bond dimension calculations.
    • These hyperparameters were optimized through a grid search strategy, ensuring that the selected values provided a good balance between performance, stability, and computational cost. In particular, we found that a bond dimension of around 120 offered sufficient expressivity without excessive memory consumption.
  4. 4.
    Anomaly Scoring:
    • For each event in the dataset (including the test data), the MPS model computes a Negative Log Likelihood (NLL) score, representing the likelihood that the event belongs to the baseline normal behavior.
  5. 5.
    Threshold Setting for Anomalies:
    • Anomalies are flagged based on a threshold applied to the NLL scores.
    • This threshold can be computed using statistical measures such as the mean and standard deviation of NLL values in the training dataset.
  6. 6.
    Evaluation:
    • The remaining 30% of the data, which includes both benign and attack-related events, is used to evaluate the model.
    • Performance metrics include detection rates for attack events and false positive rates for benign events.
@@
Our attack-detection system identifies attacks as anomalies. Notably, not all detected anomalies will be attacks. Indeed, there may be low-probability events that are perfectly legitimate (e.g., energy outages, system reboots). Because of this, it is important to adjust the parameters available in the model to minimize the detection of false positives, i.e., anomalies that are not attacks after all. With this in mind, central to the model’s success is the implementation of an acceptance threshold for the Negative Log Likelihood (NLL) of each event being generated by the model, so that events that exceed the threshold are classified as anomalous and may be considered by a human analyst to check if they correspond to a real attack. Establishing this threshold is challenging and varies in practice; some analysts may use statistical metrics like the mean and standard deviation to discern attacks within anomalies, while others may opt for the median. Notably, the average NLL of daily benign events is consistently lower than that of attacks, confirming the model’s adeptness at flagging attacks as anomalies.
Through rigorous training, which leverages 70% of the data and respects the chronological order of events, our model effectively captures attack patterns. This temporal fidelity is crucial as shuffling the order of events has been shown to influence model training negatively. The remaining 30%, containing both benign events and all attack-related events, formed the testing dataset. This approach ensures that the training data simulates realistic scenarios without prior knowledge of future events, aligning with the unsupervised nature of the MPS model. Here, the goal is to detect deviations from the baseline (training data) without explicit attack labels. This setup rigorously evaluates the model’s anomaly detection capability on unseen data, including the full spectrum of attack patterns present in the test set.
@@
In Fig.3 we show one way to determine the threshold of NLL in our model, by analyzing the number of anomalies detected and the number of attack events present in those detected anomalies for multiple threshold values (notice that, in the figure, the scales for both quantities are different). Depending on the specific goal, one can choose a low threshold value and maximize the number of attack events at the expense of having a higher false positive rate, or choose a high value for the threshold and minimize the false positive rate at the expense of not detecting all the attack events inside the detected anomalies. In any case, in the figure we can see that the number of detected anomalies decreases very quickly with the threshold, whereas this is not the case for the number of detected attacks, meaning that our MPS model is remarkably good at pinpointing anomalies that correspond to true attacks.
@@
The MPS framework does not assume independence among input features, is specifically designed to capture correlations between features through its tensor network structure. Each tensor in the MPS represents a feature, and the connections (bonds) between tensors model the interactions between features.
The bond dimension, denoted as $D$, is a critical parameter that determines the capacity of the MPS to capture feature correlations. A higher bond dimension allows the MPS to represent more complex and stronger correlations among features. Specifically, the expressiveness of the MPS in modeling joint probability distributions increases with the bond dimension, enabling it to capture higher-order dependencies that are common in cybersecurity data.
While a higher bond dimension enables the MPS model to capture richer feature correlations and detect more complex anomalies, it comes with increased computational cost. As mentioned above, the training complexity of an MPS model grows as $O(NpD^{3})$. A higher D allows for better expressivity but results in increased memory consumption and longer training times. Therefore, an optimal trade-off is needed: too low a bond dimension may cause underfitting (missing important correlations), whereas too high a bond dimension may lead to overfitting or impractical training times. Our experiments indicate that a bond dimension in the range of 50 to 150 provides a balance between performance and computational efficiency, as evidenced by the decrease in the false positive rate shown in Fig. 4. These FPR values are achieved while maintaining a high detection precision of 92.3%.
In practice, our MPS model achieves remarkable results: we see that by fine-tuning the threshold, an average of 83.5% of incident-report events are correctly identified as anomalies, and all attack types are detected (100% attack detection rate). Moreover, the average false positive rate (FPR) stands at a manageable 1.39%, substantially narrowing the search space for analysts by 98.61%.
It is worth noting that the time complexity of fitting the MPS model using the DMRG optimization process is higher than that of neural-network-based methods such as autoencoders. This is because we selected a high bond dimension to ensure top performance and interpretable insights, prioritizing explainability and accuracy over training speed. While autoencoders are simpler, faster to train, and effective in anomaly detection, they function as black-box models and lack inherent interpretability. In contrast, the MPS model offers rich insights into feature correlations and anomaly characteristics, making it a better fit for scenarios where interpretability is critical.
Notably, the results could be improved even further by increasing the MPS bond dimension, and/or by using more complex TN structures, able to capture more subtle correlations in the model. Last but not least, our MPS model can be scaled easily to datasets with a huge number of features, and is remarkably memory-efficient, on top of being accurate (as shown) and explainable. This last property is what we discuss in detail in the next section.
@@ IV.1 Direct probability extraction
As presented in subsection II.4, the RDM’s diagonal entries provide a probabilistic view of specific subsystem’s state, which is directly connected to the dataset’s frequency distribution. The MPS captures and adapts to the correlations in the data during training, allowing security analysts to assess how individual feature distributions contribute to detecting anomalous behaviors. By comparing empirical frequency distributions with MPS-derived probabilities, analysts can pinpoint deviations that may indicate cybersecurity threats. While a simple frequency distribution is akin to a static snapshot, capturing conditional probabilities at a fixed time step, MPS allows for an efficient dynamical representation of other valuable insights, turning exhaustive analysis such as the case of all-to-all feature correlations (scaling exponentially in computational complexity) into a polynomial complexity scaling.
@@ IV.2 Von Neumann entropy
The concept of Von Neumann Entropy is analogous to the Shannon entropy in classical information theory [19] but adapted for quantum systems and reduced density matrices. In the end, the Von Neuamn entropy of a RDM is nothing but the Shannon entropy of its eigenvalues. It quantifies the correlation between a subsystem with the rest of the system, thus providing insight into how independent a feature is within the broader data context. This metric plays a crucial role in cybersecurity anomaly detection, as it helps distinguish between features that contribute independently to an event and those that are contextually bound to cyber threats. Features with high entropy often indicate complex dependencies, such as coordinated attacks or lateral movements within a compromised network, while low-entropy features may correspond to isolated system events. This allows to differentiate between what might be deemed normal and anomalous within a dataset, effectively utilizing “context” to pinpoint deviations.
@@ IV.3 Feature importance
To further analyze the differences between benign and attack events, we computed the mean probabilities of feature values for both benign and attack instances as given by the MPS model. In our model, the likelihood of each event $\mathbf{v}=(v_{1},v_{2},\dots,v_{N})$ is computed as the squared magnitude of the wavefunction $\Psi(\mathbf{v})$, which, for an MPS model, can be expressed as a product of probabilities associated with each feature:
$P(\mathbf{v})=|\Psi(\mathbf{v})|^{2}=\prod_{i=1}^{N}P_{i}(v_{i})$ (10)
where $P_{i}(v_{i})$ is the probability associated with the $i$-th feature value $v_{i}$ derived from the local density matrices in the MPS model. The overall likelihood of an event is thus the product of the probabilities of its feature values. We compute the mean of these overall likelihoods separately for benign events and attack events to compare their distributions. Additionally, the Negative Log Likelihood (NLL) for an event is calculated as:
$\mathrm{NLL}=-\log P(\mathbf{v})=-\sum_{i=1}^{N}\log P_{i}(v_{i}).$ (11)
This NLL represents how unlikely an event is under the model trained on benign data; higher NLL values indicate greater anomalies.
Feature Mean Prob. (Benign) Mean Prob. (Attack)
@@
Mean Overall Likelihood 2.14e-10 6.06e-18
Mean NLL 9.669 17.217
Table 2: Mean probabilities of feature values and overall likelihoods for benign and attack instances.
Table 2 presents the computed mean probabilities for features in benign and attack contexts, which can be instrumental in feature selection. Features with high probabilities in the benign category and low probabilities in the attack category, such as ‘data.parent_process_path‘, ‘data.process_user_name‘, and ‘data.parent_process_basename‘ (features 2, 3, and 4), are indicative of normal operational patterns. Their high likelihood under benign conditions and rarity during attacks signify their importance in identifying abnormal behavior. These features represent critical aspects of process-level activity: the execution path of a parent process, the user responsible for the process, and the process’s basename. Their behavior aligns with real-world expectations, where deviations in these parameters often signal potential malicious activities, such as unauthorized user access or unusual process execution paths. Conversely, features with a higher mean in the attack category than in the benign may be less informative for detecting anomalies, as their higher occurrence in attack scenarios does not contribute to distinguishing attacks from benign behavior. In addition, the total probability values, obtained by the product of probabilities across all features, reveal a striking disparity: the probability associated with attack features is several orders of magnitude lower than that of benign features, and the mean NLL for attack events is significantly higher. This stark contrast explains why the MPS model effectively discriminates between benign operations and cyberattacks: the improbability of the conjunction of attack features underscores their anomaly.
@@ IV.6 Mutual Information
The incorporation of mutual information into our MPS framework allows to understand the intricate correlations between features within the system. The mutual information heatmap in Fig.9 is particularly valuable in cybersecurity, as it reveals which features share the most information, highlighting patterns such as correlated network events or coordinated attacker behavior. By clustering highly correlated features, security teams can optimize intrusion detection rules and reduce redundant monitoring of less informative variables. Beyond interpretability, this metric holds practical utility in optimizing the MPS model. By rearranging columns to position highly correlated features adjacent to each other, we can refine the efficiency of the virtual bond dimension, ensuring a more resource-effective and potent representation of the data’s entangled nature. In addition, those pairs of features with large mutual information could even be clustered together, therefore reducing the complexity of the problem.
@@ V Conclusions and Next Steps
In this paper we have discussed in detail the potential of TN methods to develop explainable and interpretable machine learning algorithms. We have shown the validity of the approach for an MPS algorithm designed for clustering, with the application to a real-life use case in cybersecurity. Our research has unveiled the potential of MPS as a powerful tool for unsupervised anomaly detection, particularly within the domain of adversary-generated threat intelligence. The key findings demonstrate that MPS not only rivals but in certain respects surpasses the performance of traditional machine learning models such as autoencoders and GANs, especially in terms of explainability and interpretability. The application of MPS to a cybersecurity dataset revealed its robustness in identifying cyberthreats with a low false positive rate, a critical parameter in cybersecurity operations. The ability of MPS to provide direct probability extraction, assess feature importance, and elucidate on the reasons behind the classification of data points as anomalies underscores its utility in operational environments where understanding the ‘why’ behind a model’s decision is as vital as the decision itself. In addition, the research has highlighted the interpretability capabilities of MPS, offering insights into Von Neumann entropy, feature importance, conditional probabilities, and mutual information directly from the model’s parameters without the need for additional tools or techniques. These capabilities have profound implications for explainable AI, setting a new benchmark for transparency in machine learning applications. In addition, the methods presented in this paper allow the capability to generate synthetic data in order to improve training of complex models (due to lack of data availability) and activity generation for deception systems. This would also help in the development of behavior-based models, which are better against unknown attacks than rule-based models.
The methods developed in this work can be improved in many ways. For instance, we could consider more advanced TN structures beyond MPS, with more intricate patterns of correlations. Furthermore, as part of future work, we aim to validate our MPS-based approach on additional well-established cybersecurity datasets to assess its robustness across different data distributions. Specifically, we plan to apply our model to:
  • KDD Cup 1999 [21]: One of the most widely used intrusion detection datasets, containing simulated attacks in a network traffic environment.
  • CIC-IDS2017 [22]: A modern intrusion detection dataset with real-world traffic and attack types collected by the Canadian Institute for Cybersecurity.
  • UNSW-NB15 [23]: A dataset capturing hybrid real and synthetic network traffic to simulate real-world cyber threats.
These datasets contain a diverse set of attack types and real-world traffic patterns, making them ideal benchmarks for evaluating the scalability and robustness of our method. This future validation will allow us to assess whether our model retains its effectiveness across varying data distributions and attack scenarios. We leave this as an important direction for future research.
@@
Acknowledgements.- We acknowledge Donostia International Physics Center (DIPC), Ikerbasque, Basque Government, Diputación de Gipuzkoa, European Innovation Council (EIC) and Tecnun for constant support, as well as insightful discussions with the teams from Multiverse Computing, DIPC and Tecnun on the algorithms and technical implementations. Special thanks to the team at CounterCraft SL, for providing the motivation for this work, the dataset and numerous technical discussions. This work was supported by the Diputación Foral de Gipuzkoa through the “Q4DecOps: Algoritmo cuántico para detección de anomalías en entornos de ciber-engaño” project (2022-QUAN-000026-01).